OpenStudy (anonymous):

Given the graphs of the linear function, f(x) and the quadratic g(x), evaluate: (f+g)(2)...

OpenStudy (anonymous):

this is the graphhhh

OpenStudy (anonymous):

could it be any harder to read?

OpenStudy (anonymous):

looks kinda like \(f(2)=2\) right?

OpenStudy (anonymous):

and maybe \(g(2)=-4\) if i am reading it right

OpenStudy (anonymous):

like i said, kinda hard to read but if that is right then \[(f+g)(2)=f(2)+g(2)=2+(-4)=-2\]

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

you got this?

OpenStudy (anonymous):

@satellite73 yeah no i understood what i needed to do i just needed clarification. thank youuu